1. How many hosts can reside on a single Class C network?

1.

65535

2.

256

3.

254

4.

127

2.  

Select the invalid IP address

1.

95.31.0.15

2.

143.256.11.2

3.

24.123.33.60

4.

135.10.243.1

3.

What layer in the TCP/IP protocol suite would a Web Browser reside?

1.

Data Link

2.

Application

3.

Network

4.

Physical

4.

If I have developed a new type of network interface device - one that uses a new wireless technology, which layer of the TCP/IP protocol suite would I need to modify to support it?

1.

IP

2.

Application

3.

TCP

4.

Data Link

5.

What is the name of the service that converts something like www.myinternetsite.com to the IP address of the computer?

1.

Address Resolution Protocol

2.

Reverse Address Resolution Protocol

3.

Domain Name Service

4.

Router
Answer #1

ANSWERS

QUESTION 1. 254

QUESTION 2.  143.256.11.2

QUESTION 3. Application

QUESTION 4. Data Link

Explanation
1. 2^8 - 2 = 256 - 2 = 254
2. Range for each number is 0-255 since there are only 8 bits. So, 256 is an invalid entry
3. Web browser is an application to view only. So, it is under Application layer
4. Data link since it makes the link contact

-- 4 sub parts at a time please - Policy of Chegg